OpenStudy (anonymous):

use implicit differentiation to find the slope of the tangent line at (2,1) and use it to find the equation of the tangent line in y=mx + b form

OpenStudy (anonymous):

heres the equation \[3^x+\log_2_(xy)=10\]

OpenStudy (anonymous):

ok sorry one more time \[3^x+\log_2(xy)=10\] thats the right equation

OpenStudy (anonymous):

and that is log base 2 if you can't read that

OpenStudy (anonymous):

\[( Ln2*2^{x} + 1/xyLna *y) + ( 1/xyLna *x)y' = 0\] that would be the implicit derivative

OpenStudy (anonymous):

i have no clue how you did that haha

OpenStudy (anonymous):

now: express y' from this expression and you got the slope

OpenStudy (anonymous):

take partial derivatives, :)

OpenStudy (anonymous):

i hope i didnt make any mistakes in derivation, make sure by ur self

OpenStudy (anonymous):

\[d/dx[a ^{x}] =Ln(a)*a ^{x}\]

OpenStudy (anonymous):

the rest of derivatives are simple

OpenStudy (anonymous):

\[ d/dx[\log_{a} x]=1/xLn(a)\]

OpenStudy (anonymous):

got it?

OpenStudy (phi):

Can you do the differentiation? I would change \( 3^x \) to \( e^{x ln3} \) its derivative is \( ln 3 e^{x ln3} = ln(3) 3^{x}\)

OpenStudy (phi):

rewrite log_2(xy) as \[ \frac{1}{ln(2)} ln(xy) \] its derivative is \[ \frac{1}{ln(2)} \frac{1}{xy} (x dy/dx + y dx/dx) \] or just \[ \frac{1}{ln(2)} \frac{1}{xy} (x \frac{dy}{dx} + y ) \]

OpenStudy (phi):

multiply through by ln(2)xy: \[ y ln(2)ln(3) x 3^x + y + x \frac{dy}{dx} = 0 \] and \[ \frac{dy}{dx} =-\frac{y}{x}(ln(2)ln(3)x3^x+1) \]